# Spatially resolved broad line region in a quasar at z=4

### v1

@@ 3.1.1 BLR size from scaling relations
1.74 | Bentz et al. (2013) | commonly used |
1.47 | Du et al. (2018) | lower subsample |
1.22 | Du & Wang (2019) | without correction |
0.94 | Du & Wang (2019) | with correction |
0.65 | Woo et al. (2024) | 47 AGN sample |
0.50 | Du et al. (2018) | higher subsample |
0.48 | Woo et al. (2024) | 240 AGN sample |
We have derived the BLR size from 4 different relations and summarised them in Table 2. One of the best known relations between the AGN luminosity and size of the BLR is that of Bentz et al. (2013). The AGN luminosities were derived using a surface brightness decomposition to correct for the host galaxy; and the BLR sizes were derived from a simple measure of the time lag obtained from RM. An emphasis to AGN with high accretion rates was given by Du et al. (2018), who derived different relations depending on the accretion rate. We show here the relation for both lower and higher accretion rates, which yield very different results. We also consider a version of that relation which was updated by Du & Wang (2019) to include an explicit correction for the accretion rate via the parameter. For J0529 this provides a substantial reduction to the inferred BLR size compared to Bentz et al. (2013), although for J0529 the correction itself is only a 25% effect. Finally, we consider estimates from Woo et al. (2024) who provided two fits. We include both, one based on a combination of 47 AGN from two specific samples, and one derived from 240 AGN that are mostly from different sources in the literature. These authors found no strong dependent deviation from a relation, but they did note that the slope of the relation is likely to be shallower than that popularly used. These differences highlight the uncertainties associated with employing scaling relations. Since there is also a scatter (typically reported as dex) associated with each relation, here we can conclude only that is likely in the range 0.5-1.7 pc.
@@
10.92 | Dalla Bontà et al. (2020) | Eq. 38 using & LHβ |
10.24 | Wolf et al. (2024) | continuum modelling; also C IV & Mg II |
10.16 | Woo et al. (2015) | Eq. 2 using & L5100 |
9.87 | Reines et al. (2013) | Eqs. 2, 3, & 5; converting FWHMHβ & L5100 to H |
9.58 | Vietri et al. (2020); Bongiorno et al. (2014) | Eq. 4 or Eq. 2 respectively, using FWHMHβ & L5100 |
3.1.2 Black hole mass from scaling relations
@@
In a similar approach, we estimate the black hole mass from several different relations which are summarised in Table 3. The mass reported by Wolf et al. (2024) in their discovery paper was based on both modelling of the continuum shape and use of the C IV 1549 and Mg II 2799 line widths. A recent re-assessment of the use of H to estimate was provided in Eq. 38 of Dalla Bontà et al. (2020), who included a (statistical) correction for Eddington ratio . These authors also argued that the dispersion is a less biassed tracer of the line width than FWHM (although use of for profiles with Lorentzian-like wings also poses difficulties, Santos et al. 2025); and prefer to use the line rather than continuum luminosity in order to avoid bias from core dominated radio sources. An important part of the relation is converting the virial mass into a black hole mass via the parameter, and Dalla Bontà et al. (2020) adopt from Batiste et al. (2017). For J0529, the inferred BH mass is . A likely reason for such a high mass is the bias on both the line dispersion and luminosity caused by the blue wing. Indeed, using only the Gaussian corresponding to the main part of the line profile yields a lower mass consistent with that of Wolf et al. (2024). An earlier relation which did not attempt to make any corrections, and was based on the relation of Bentz et al. (2013), is given as Eq. 2 of Woo et al. (2015), for which they adopt (a very different value than used by Dalla Bontà et al. (2020), emphasizing that it depends on how the scaling relation is set up and whether FWHM or is used). From this relation we infer a value of using the full line profile. Reines et al. (2013) derived a relation based on the continuum luminosity and FWHM of the H line, providing conversions for these quantities to H luminosity and FWHM in Eqs. 2 and 3. This also builds on the relation of Bentz et al. (2013), and uses the conversions to H from Greene & Ho (2005). We can put our measured values into those equations, apply the conversions, and hence estimate a mass from their Eq. 5. Doing this, and noting that they use , yields . The final relation we consider is Eq. 4 from Vietri et al. (2020) which was applied to luminous QSOs at , and which was taken directly from Eq. 2 in Bongiorno et al. (2014) where it was applied to X-ray obscured red QSOs at . As for the previous relation it uses FWHMHβ and L5100, yielding . The conclusion here is that is associated with even more uncertainty than , and that there may be a systematic difference when using vs FWHM. For J0529 the scaling relations tend to give values around .
3.2 Comparison to UV lines
In this section we compare the H I line profile to the UV lines C IV and O I in the Xshooter spectrum published by Wolf et al. (2024). To do so, we have subtracted an estimate of the true continuum level as indicated by the lowest regions in the spectral ranges around the two lines. We have not tried to fit the iron emission, nor other neighbouring emission, since the aim is only to compare the line profiles.
The right panel of Fig. 3 shows the comparison to O I 1302 Å, a low excitation line that in AGN is expected to arise in the BLR through Ly fluorescence (Netzer & Davidson, 1979; Kwan & Krolik, 1981; Matsuoka et al., 2007). O I 8446 Å is also enhanced through this process and the ratio between these lines in the resulting radiative cascade can provide a measure of the extinction towards the BLR (Netzer & Davidson, 1979). Given that the Balmer line ratios in Sec. 3 hint that there might be some modest extinction, an independent measure like this would enable one to clarify the impact. There is a good match on the red side of the profiles, but the O I line shows no evidence for a blue wing suggesting it is more likely associated with the systemic velocity of the QSO. We note that the peak of the O I line is contaminated by deep absorption across a range of velocities, the cause of which is unclear.
The C IV 1549 Å line in the left panel presents a contrasting picture since it is only associated with the outflow, and can be traced to a velocity of . It confirms the nuclear nature of the outflow, and from the size-luminosity relation of the C IV line and 1350 Å continuum (based on 7-13 yr RM campaigns) we infer a characteristic radius for the C IV emission of 0.4 pc (Kaspi et al., 2021; Shen et al., 2024). Analysis of the C IV line is not our main aim, and it is often done in association with [O III] emission which is outside our spectral range. We therefore note only that the line properties are often grouped into two classes of QSOs related to whether the C IV is dominated by a more systemic or strongly blue-shifted component, which are also broadly linked to the properties of Narrow Line Seyfert 1 galaxies (Sulentic et al., 2011; Sulentic & Marziani, 2015; Marziani et al., 2017; Vietri et al., 2018, 2020). In this scheme the moderate equivalent width, blue-shifted C IV line in J0529 would be associated with a disk wind.
Because the C IV line is easily observable in high redshift QSOs and is known to originate in the BLR, it has been used to estimate despite the often unusual line properties (Vestergaard & Peterson, 2006; Park et al., 2013; Runnoe et al., 2013). While this is physically reasonable for QSOs in which the line is both approximately symmetric and close to the systemic velocity, it is harder to explain a relation between the line properties and for emission originating in an outflow. Despite this, Coatman et al. (2017) found an empirical relation between the line’s blue-shift and a correction factor that can be applied to the inferred mass, to match it to other scaling relations. However, Temple et al. (2023) have shown that the situation is more complex, and that there is a 2-dimensional parameter space relating the blue-shift (and equivalent width) to a combination of and . They showed there is a wedge in that parameter space where the C IV blue-shift increases (and equivalent width decreases), and that in general the highest blue-shifts are observed only in AGN with both and . Their modelling to reproduce these effects is bringing us to a better understanding of the cause of the two populations of QSOs with differing line properties noted above. J0529 has a blue-shift km -1 higher than their sample range, and has both a high and , consistent with their analysis.
@@ 4 Description of the BLR model
We make use of the phenomenological model described by Pancoast et al. (2014), our implementation of which is given in detail in GRAVITY Collaboration et al. (2020) with summaries in GRAVITY Collaboration et al. (2021) and GRAVITY Collaboration et al. (2024). The model implements the BLR as an ensemble of clouds moving in the potential of the central black hole. Physically, clouds in the BLR are likely to be transient objects, plausibly formed by shock compression due to gas collisions on small scales (as has been adopted to model the ISM on scales of -10 pc in AGN, Vollmer et al. 2018). However, irrespective of the true nature of the clouds, in our phenomenological model the use of the term is for convenience. The clouds should be taken to represent line emitting test particles rather than physical entities, because the model parameterizes the distribution and kinematics of the line emission rather than the entire gas content of the BLR.
@@
The model of Pancoast et al. (2014) is a good representation of the physics motivating the FRADO (Failed Radiatively Accelerated Disk Outflow) model, proposed by Czerny & Hryniewicz (2011) and developed by Czerny et al. (2017) and Baskin & Laor (2018). In this concept, the anisotropic emission from the accretion disk and the effects of shielding enable dust to exist in a thin disk at radii smaller than the nominal sublimation radius. The radiation pressure from the hot disk can accelerate a dusty gas cloud vertically until it has a direct line of sight to the accretion disk. At this point the dust sublimates, and the cloud continues ballistically before falling back down to the disk. The BLR therefore forms a flared region above the disk plane, as proposed by other authors (Goad et al., 2012; Ramolla et al., 2018; Hopkins, 2025). These models are consistent with the dust having a bowl-shaped form that was proposed to explain the difference in dust sizes between interferometric and reverberation mapping measurements (Gravity Collaboration et al., 2024).
@@
A limitation of the FRADO model is that it only considers vertical motion of the clumps, and does not address their radial motion. Once clouds are driven high enough that the dust sublimates, the direct line of the sight they have to the AGN is also expected to lead to an outward acceleration. This may be either by UV-line driving, as is believed to be happening for ultra fast outflows (UFOs) from the accretion disk (Mizumoto et al., 2021), or as a magneto-centrifugal wind (Emmering et al., 1992), which has also been proposed as an outflow mechanism on 1-10 pc scales (Vollmer et al., 2018). The BLR outflow will form part of a hierarchy of outflows, located between the small scale UFOs and the radiatively driven outflow that accelerates dusty gas to create geometrically thick obscuration (Ramos Almeida & Ricci, 2017; Chan & Krolik, 2017; Hönig, 2019). Physically, and based on simulations (Kudoh et al., 2023, 2024; Hopkins et al., 2025; Hopkins, 2025), one would expect outflow to be driven at the surface of the BLR and inflow to occur closer to the disk plane. However, in the Pancoast et al. (2014) model, radial motions are included in a way that allows a fraction of the clouds throughout the bulk of the BLR to be either all inflowing or all outflowing (and unbound) on trajectories that are essentially radial.
@@
In order to better physically motivate the way radial flows are handled in our phenomenological model, we have updated its implementation for the analysis here, while keeping the model as simple as possible. The original flared disk model is used to represent the static part of the BLR described by Czerny et al. (2017) and Baskin & Laor (2018). It is unchanged except that outflows are excluded, while inflow is in principle allowed. However, after initial iterations of the model, it was clear that for J0529 the fraction of inflowing clouds was small enough that, for simplicity, it could be excluded entirely with negligible impact. Thus, in this case, the motions in the disky part of the model are purely rotational.
@@
A specific outflow component is added as an additional wedge that extends above the upper surface of that flared disk as shown in Fig. 4. Again for simplicity, to avoid including more free parameters, we assume that the outflow and the disk components are spatially connected. It represents the concept put forward by Elvis (2000) as well as the outflowing part in the simulations of Kudoh et al. (2023, 2024) and Hopkins et al. (2025), and has been modelled explicitly by Yong et al. (2017, 2020). The total angular height of the BLR is then the sum of for the flared disk and for the outflow. Thus a situation where would mean that there are clouds on orbits perpendicular to the disk plane.
@@
Clouds in the outflow are required to be rotating around the central black hole with the same circular velocity they would have if they were in the midplane at the same projected radius. They must also be outflowing with a radial velocity . The original model defined the outflow velocity at a radius to be the same as the escape velocity at that radius. A couple of examples show that this was a reasonable assumption. For and derived in NGC 3783 (GRAVITY Collaboration et al., 2021), where about half the clouds are on unbound orbits, km s-1 at the average radius of the BLR is only a factor greater more than the 1800 km s-1 FWHM and within the full line profile; Similarly, for the nominal size (1 pc) and mass ( M⊙) inferred in Sec. 3.1 for J0529, V km s-1 is only a factor of a few greater than the dispersion of the line profile including the blue wing. However, the assumption that is also a limitation, and can lead to unphysical geometries for prominent fast outflows for which the model is not suited. Rather than follow the same prescription for we apply an improvement suggested by Pancoast et al. (2014) to link it to the escape velocity at the wind launching radius (Murray et al., 1995; Long et al., 2023). Our implementation ties Vout more flexibly to the escape velocity at the minimum radius and still enables it to decrease with radius:
@@
where is the minimum radius for cloud distribution, is a scaling factor (expected to be in the range one to a few) for the escape velocity at , and is a negative index in the range allowing to decrease with increasing radius. We argue that such a radial decrease should be expected because reverberation mapping of the C IV line, which has much a higher excitation potential than H I, tends to yield a smaller radial scale than H (Lira et al., 2018; Shen et al., 2024); and the line is often blue-shifted with respect to the H I lines (Coatman et al., 2017; Temple et al., 2023), as also illustrated for J0529 in Fig. 3. Thus it is reasonable to conclude that the outflow velocity is higher at smaller radii.
@@ 5 The fitted BLR model
We fit the H and H lines simultaneously in a similar way to that described in Kuhn et al. (2024): all the parameters in the fit are tied to be the same for both lines except for the two that define their radial distribution. These are the parameter that defines the form of the distribution, and the scale factor which is linked to the average radius such that , where is the same minimum radius for both lines. The details of the fit are given in Appendix A, including Table 4 summarising the best fitting parameters corresponding to the model, and Fig. 8 of the resulting spectra and differential phases for each baseline. As a complement to this, the top two rows of Fig. 5 show the data and fitted model for each of the line profiles as well as the sum of the phases from baselines UT 4-2, 4-1, and 3-2 which contain the strongest signal. The three main points from the phase data in Fig. 5 are (i) the phase signal for H is about 1/3 the strength of that for the H line, as expected from their equivalent widths; (ii) the phase signal is single-sided which is indicative of predominant outflow rather than the S-shape associated with rotation; and (iii) there are two distinct peaks in the phase signal, which are related to the separate outflowing and disky components. This last point is developed further in the bottom panels of Fig. 5 which show the contribution of the outflowing and disky components to the total line profile. It is clear that emission from the outflow is not just limited to the blue wing of the profile, but makes up a substantial part of the core as well. In contrast the disky component has only a subdominant contribution on the red side of the line. As such, the model implies that the systemic velocity of J0529, which should be traced by the disky component of the BLR that peaks at 2.423 m, is offset by about 1200 km s-1 from the peak of the line. This would indicate an update of the redshift to .
Histograms of the posterior distributions for all the parameters are presented in Fig. 9, and the most relevant results for our analysis are additionally shown in the corner plot in Fig. 10. These figures indicate that the fit is well constrained; and that while other potential solutions might exist, these are at a much lower probability.

### v2

@@ 3.1 BLR size from scaling relations
| 1.74 | ben13 | commonly used |
| 1.47 | du18 | lower subsample |
| 1.22 | du19 | without correction |
| 0.94 | du19 | with correction |
| 0.65 | woo24 | 47 AGN sample |
| 0.50 | du18 | higher subsample |
| 0.48 | woo24 | 240 AGN sample |
We have derived the BLR size from 4 different relations and summarised them in Table 2. One of the best known relations between the AGN luminosity and size of the BLR is that of ben13. The AGN luminosities were derived using a surface brightness decomposition to correct for the host galaxy; and the BLR sizes were derived from a simple measure of the time lag obtained from RM. An emphasis to AGN with high accretion rates was given by du18, who derived different relations depending on the accretion rate. We show here the relation for both lower and higher accretion rates, which yield very different results. We also consider a version of that relation which was updated by du19 to include an explicit correction for the accretion rate via the parameter. For J0529 this provides a substantial reduction to the inferred BLR size compared to ben13, although for J0529 the correction itself is only a 25% effect. Finally, we consider estimates from woo24 who provided two fits. We include both, one based on a combination of 47 AGN from two specific samples, and one derived from 240 AGN that are mostly from different sources in the literature. These authors found no strong dependent deviation from a relation, but they did note that the slope of the relation is likely to be shallower than that popularly used. These differences highlight the uncertainties associated with employing scaling relations. Since there is also a scatter (typically reported as dex) associated with each relation, here we can conclude only that is likely in the range 0.5-1.7 pc.
@@
| 10.43 | dal20 | Eq. 38 using & LHβ |
| 10.24 | wol24 | continuum modelling; also C IV & Mg II |
| 10.16 | woo15 | Eq. 2 using & L5100 |
| 9.87 | rei13 | Eqs. 2, 3, & 5; converting FWHMHβ & L5100 to H |
| 9.58 | vie20; bon14 | Eq. 4 or Eq. 2 respectively, using FWHMHβ & L5100 |
3.2 Black hole mass from scaling relations
@@
In a similar approach, we estimate the black hole mass from several different relations which are summarised in Table 3. The mass reported by wol24 in their discovery paper was based on both modelling of the continuum shape and use of the C IV 1549 and Mg II 2799 line widths. A recent re-assessment of the use of H to estimate was provided in Eq. 38 of dal20, who included a (statistical) correction for Eddington ratio . These authors also argued that the dispersion is a less biassed tracer of the line width than FWHM (although use of for profiles with Lorentzian-like wings also poses difficulties, san25); and prefer to use the line rather than continuum luminosity in order to avoid bias from core dominated radio sources. An important part of the relation is converting the virial mass into a black hole mass via the parameter, and dal20 adopt from bat17. For J0529, the inferred BH mass is , which is high probably because of the bias on both the line dispersion and luminosity caused by the blue wing. An earlier relation which did not attempt to make any corrections, and was based on the relation of ben13, is given as Eq. 2 of woo15, for which they adopt (a very different value than used by dal20, emphasizing that it depends on how the scaling relation is set up and whether FWHM or is used). From this relation we infer a value of using the full line profile. rei13 derived a relation based on the continuum luminosity and FWHM of the H line, providing conversions for these quantities to H luminosity and FWHM in Eqs. 2 and 3. This also builds on the relation of ben13, and uses the conversions to H from gre05. We can put our measured values into those equations, apply the conversions, and hence estimate a mass from their Eq. 5. Doing this, and noting that they use , yields . The final relation we consider is Eq. 4 from vie20 which was applied to luminous QSOs at , and which was taken directly from Eq. 2 in bon14 where it was applied to X-ray obscured red QSOs at . As for the previous relation it uses FWHMHβ and L5100, yielding . The conclusion here is that is associated with even more uncertainty than , and that there may be a systematic difference when using vs FWHM. For J0529 the scaling relations tend to give values around .
@@ 4 Description of the BLR model
We make use of the phenomenological model described by pan14, our implementation of which is given in detail in gra20sha with summaries in gra21shi and gra24san. The model implements the BLR as an ensemble of clouds moving in the potential of the central black hole. Physically, clouds in the BLR are likely to be transient objects, plausibly formed by shock compression due to gas collisions on small scales (as has been adopted to model the ISM on scales of -10 pc in AGN, vol18). However, irrespective of the true nature of the clouds, in our phenomenological model the use of the term is for convenience. The clouds should be taken to represent line emitting test particles rather than physical entities, because the model parametrises the distribution and kinematics of the line emission rather than the entire gas content of the BLR.
@@
The model of pan14 is a good representation of the physics motivating the FRADO (Failed Radiatively Accelerated Disk Outflow) model, proposed by cze11 and developed by cze17 and bas18. In this concept, the anisotropic emission from the accretion disk and the effects of shielding enable dust to exist in a thin disk at radii smaller than the nominal sublimation radius. The radiation pressure from the hot disk can accelerate a dusty gas cloud vertically until it has a direct line of sight to the accretion disk. At this point the dust sublimates, and the cloud continues ballistically before falling back down to the disk. The BLR therefore forms a flared region above the disk plane, as proposed by other authors (goa12; ram18; hop25b). These models are consistent with the dust having a bowl-shaped form that was proposed to explain the difference in dust sizes between interferometric and reverberation mapping measurements (gra24cao).
@@
A limitation of the FRADO model is that it only considers vertical motion of the clumps, and does not address their radial motion. Once clouds are driven high enough that the dust sublimates, the direct line of the sight they have to the AGN is also expected to lead to an outward acceleration. This may be either by UV-line driving, as is believed to be happening for ultra fast outflows (UFOs) from the accretion disk (miz21), or as a magneto-centrifugal wind (emm92), which has also been proposed as an outflow mechanism on 1-10 pc scales (vol18). The BLR outflow will form part of a hierarchy of outflows, located between the small scale UFOs and the radiatively driven outflow that accelerates dusty gas to create geometrically thick obscuration (ram17; cha17; hoe19). Physically, and based on simulations (kud23; kud24; hop25a; hop25b), one would expect outflow to be driven at the surface of the BLR and inflow to occur closer to the disk plane. However, in the pan14 model, radial motions are included in a way that allows a fraction of the clouds throughout the bulk of the BLR to be either all inflowing or all outflowing (and unbound) on trajectories that are essentially radial.
@@
In order to better physically motivate the way radial flows are handled in our phenomenological model, we have updated its implementation while keeping the model as simple as possible. The original flared disk model is used to represent the static part of the BLR described by cze17 and bas18. It is unchanged except that outflows are excluded, while inflow is in principle allowed. However, after initial iterations of the model, it was clear that for J0529 the fraction of inflowing clouds was small enough that, for simplicity, it could be excluded entirely with negligible impact. Thus, in this case, the motions in the disky part of the model are purely rotational.
@@
A specific outflow component is added as an additional wedge that extends above the upper surface of that flared disk as shown in Fig. 3. Again to avoid including more free parameters, we assume that the outflow and the disk components are spatially connected. It represents the concept put forward by elv00 as well as the outflowing part in the simulations of kud23; kud24 and hop25a, and has been modelled explicitly by yon17; yon20. The total angular height of the BLR is then the sum of for the flared disk and for the outflow. Thus a situation where would mean that there are clouds on orbits perpendicular to the disk plane.
@@
Clouds in the outflow are required to rotate around the central black hole with the same circular velocity they would have if they were in the midplane at the same projected radius. They are also outflowing with a radial velocity . The original model defined the outflow velocity at a radius to be the same as the escape velocity at that radius. A couple of examples show that this was a reasonable assumption. For and derived in NGC 3783 (gra21shi), where about half the clouds are on unbound orbits, km s-1 at the average radius of the BLR is only a factor greater more than the 1800 km s-1 FWHM and within the full line profile; Similarly, for the nominal size (1 pc) and mass ( M⊙) inferred in Sec. 3 for J0529, V km s-1 is only a factor of a few greater than the dispersion of the line profile including the blue wing. However, the assumption that is also a limitation, and can lead to unphysical geometries for prominent fast outflows for which the model is not suited. Rather than follow the same prescription for we apply an improvement suggested by pan14 to link it to the escape velocity at the wind launching radius (mur95; lon23). Our implementation ties Vout more flexibly to the escape velocity at the minimum radius and still enables it to decrease with radius:
@@
where is the minimum radius for the cloud distribution, is a scaling factor (in the range one to a few) for the escape velocity at , and is an index in the range allowing to decrease with increasing radius. We argue that such a radial decrease should be expected because reverberation mapping of the C IV line, which has much a higher excitation potential than H I, tends to yield a smaller radial scale than H (lir18; she24); and the line is often blue-shifted with respect to the H I lines (coa17; tem23), as also illustrated for J0529 in Fig. 6. Thus it is reasonable to conclude that the outflow velocity is higher at smaller radii.
@@ 5 The fitted BLR model
We fit the H and H lines simultaneously in a similar way to that described in kuh24: all the parameters in the fit are tied to be the same for both lines except for the two that define their radial distribution. These are the parameter that defines the form of the distribution, and the scale factor which is linked to the average radius such that , where is the same minimum radius for both lines. The details of the fit are given in Appendix A, including Table 4 summarising the best fitting parameters corresponding to the model, and Fig. 7 of the resulting spectra and differential phases for each baseline. As a complement to this, the top two rows of Fig. 4 show the data and fitted model for each of the line profiles as well as the sum of the phases from baselines UT 4-2, 4-1, and 3-2 which contain the strongest signal. The three main points from the phase data in Fig. 4 are (i) the phase signal for H is about 1/3 the strength of that for the H line, as expected from their equivalent widths; (ii) the phase signal is single-sided which is indicative of predominant outflow rather than the S-shape associated with rotation; and (iii) there are two distinct peaks in the phase signal, which are related to the separate outflowing and disky components. This last point is developed further in the bottom panels of Fig. 4 which show the contribution of the outflowing and disky components to the total line profile. It is clear that emission from the outflow is not just limited to the blue wing of the profile, but makes up a substantial part of the core as well. In contrast the disky component has only a subdominant contribution more towards the red side of the line. Histograms of the posterior distributions for all the parameters are presented in Fig. 8, and the most relevant results for our analysis are additionally shown in the corner plot in Fig. 9. These figures indicate that the fit is well constrained; and that while other potential solutions might exist, these are at a much lower probability.
5.1 Systemic velocity
Our spectrum and analysis imply a modest update of the redshift of J0529 with respect to the value of reported by wol24. Taking the peak of the H I lines as representative of the systemic velocity, we find . However, the model indicates that the systemic velocity is shifted with respect to the line peak by about 1200 km s-1. This shift is denoted by the parameter in Table 4, which is the relative shift of the zero (systemic) velocity wavelength from the initial value via the relation – hence the value of is less important than that of which it generates. The posterior distributions in Fig. 8 show that , and therefore also , is tightly constrained by the model. While there is no specific feature in the data that determines what must be, it arises because of the way the model is defined and how it is constrained by the data. The lower part of the conical outflow is aligned towards us, giving rise to the high velocities in the long blue-shifted tail where the line flux is lower because we are seeing the back side of the clouds, and there is only a weak phase signal because there is no spatial offset. In contrast, Fig. 5 shows that the upper part of the same conical outflow is viewed at a more inclined angle. Here, the lower line-of-sight velocities contribute more to the central part of the line profile as shown in Fig. 4, and the emission is stronger because we are seeing more the illuminated side of the clouds. Importantly, the spatial offset of this part of the outflow is what leads to the strong one-sided phase signal near the line centre in Fig. 4. Because of these constraints, the symmetric disk contribution must be slightly redshifted with respect to the line peak. As a test to verify this outcome, we tried modelling the data with a prior on forcing the systemic velocity to be close to the line peak. There were two peaks in the posterior distribution. The prior gave more weight to the solution close to the line peak, but this model is physically less plausible due to the high inclination and even lower black hole mass. The higher peak was nearer the edge of the prior range and had a redshifted systemic velocity, which tends to confirm the original result. This test emphasizes that the uncertainty on the systemic velocity results from matching the data – which provide strong constraints on possible geometries – within the scope of the model. Taking this uncertainty into account, the resulting redshift is . We note, however, that the difference to the previous estimates is small enough that the impact on the inferred properties of the QSO (luminosity, black hole mass, etc) is negligible.